Question 6 5100 J of heat are added to a system, it does 5900 J of work on the surroundings. What is ΔE in J? Not changed since last attempt Marked out of 1.00 Answer 11000 Flag question

11000 is wrong.

My work:

Hint: System does work on surroundings, so it loses. Wording can be tricky in these questions.

What is the correct answer then?
Answer #1

Since heat energy add to the system, hence q = 5100 J

Since work done on the surrounding it means W = - 5900 J

We know ∆E = q + w = 5100 J - 5900 J = -800 J